Suppose the yield to maturity on a two year bond is currently (3/2020) 3.2%, the current (3/2020) yield to maturity on a one year bond is 2.1%, and the term premium on a two year bond is .75%. Then, according to the liquidity premium hypothesis, the yield to maturity on a bond that matures in one year that is expected next year (3/2021) is ____%.
Answer #1

Let, required yield to maturity = R

Then,

(1+R)*(1+2.1%) = (1+(3.2% - .75%))^2

R = (1+(3.2% - .75%))^2 / (1+2.1%) - 1

R = 2.8%

So,

Yield to maturity (YTM) on a bond that matures in one year, expected next year (3/2021) is 2.8%.